OpenStudy (anonymous):

particle P satisfy the differential equation (c is a constant vector) dr/dt = c x r . Show that P moves with a constant speed on a circular path

OpenStudy (anonymous):

\[\Large \frac{d\mathbf{r}}{dt}=\mathbf{c}\times \mathbf{r}\]

OpenStudy (anonymous):

Show that the equation is satisfied by \[ \vec{r}(t) =a <\cos(\omega t), \sin(\omega t) > \] where a and omega are constants.

OpenStudy (anonymous):

is there other solution that satisfy the equation? if not, why?

OpenStudy (anonymous):

and is there any other way to solve this question without explicitly stating the components of r ?

OpenStudy (anonymous):

You can first verify that the magnitude of the velocity vector remains constant, and then that the magnitude of the position vector remains constant as well.

OpenStudy (anonymous):

give me clue how to verify that the magnitude of the velocity vector remains constant

OpenStudy (anonymous):

\[ \frac{d}{dt} v^2 = \frac{d}{dt} |\textbf{v}\cdot \textbf{v}| = 2\textbf{v}\cdot \frac{d \textbf{v}}{dt} \]

OpenStudy (anonymous):

oh wow, that worked, how do you see that coming? and last question, showing that it is circular